OpenStudy (anonymous):

Simplify the given expression. 3 over 4 i negative 3 over 4 3 I over 4 negative 3 I over 4 12 I over 16 I squared

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" what this means is to write in standard from \[\frac{3}{4i}=\frac{3}{4i}\times \frac{i}{i}=\frac{3i}{-4}=-\frac{3}{4}i\]